OpenStudy (anonymous):

Can anyone explain this to me: "An electron is accelerated from rest to 3.0 X 10^6 m/s in 5.0 X 10^-8 s. What distance does the electron travel in this time interval?" The book says 75mm, but i got 150mm. Any idea why i need to divide my answer by 2???? Please help!

OpenStudy (amistre64):

what formulas do you have to work with?

OpenStudy (amistre64):

we can work them with some calculus if need be

OpenStudy (anonymous):

For formulas i was simply multiplying the two together so as to cancelk out the s

OpenStudy (anonymous):

If there's a calculus formula, that would be great..... this is from a calc based physics class

OpenStudy (anonymous):

You probably derived it incorrectly. Did you end with d=vt where v is the final velocity?

OpenStudy (anonymous):

I don't think you need calc for this.

OpenStudy (amistre64):

if we assume a constant acceleration, then a(t) = k we are given a velocity that it reaches in a certain time limit, a boundary value. v(t) is the integration of a(t): v(t) = kt + c

OpenStudy (anonymous):

I dont think calc is really needed either. And as far as the question from the book, thats it word for word

OpenStudy (amistre64):

distance is the integration of velocity giving us: s(t) = kt^2/2 + ct + d

OpenStudy (amistre64):

since the intial velocity is at rest, c=0 when t=0 and we can determine k

OpenStudy (amistre64):

and d=0 when t=0 since we are at the start of it ...

OpenStudy (amistre64):

calc isnt needed if you have the formulas given to you already

OpenStudy (anonymous):

i kinda just thought this was a simple conversion problem....

OpenStudy (amistre64):

but i work better with the calculus way :)

OpenStudy (amistre64):

v = 3.0 X 10^6 m/s in t = 5.0 X 10^-8 s k = v/t s(t) = kt^2/2

OpenStudy (amistre64):

so: vt^2/2t = vt/2

OpenStudy (amistre64):

which is why its divided by 2 lol

OpenStudy (anonymous):

For the non-calc way,  Solve the second equation for a and then plug and chug. (Just in case you want this)